given that C=C1+cy and C1=500, c=3/4, y=1000

which one of the following statements is correct

(1) total consumption is R 500
(2) total consumption is R 1250
(3) aggregate demand is R1000
(4) marginal propensity to consume is 1/4
1
Expert's answer
2019-09-16T10:30:57-0400

(2) total consumption is R 1250

C=500+3/4*1000 = 500 + 3*1000/4 = 500 + 3*250 = 1250